National Inc. manufactures two models of CMD that can be used as cell phones, MPX, and digital camcorders.

Model Annual Sales
in Units
High F 11,000
Great P 17,000

National uses a volume-based costing system to apply factory overhead based on direct labor dollars. The unit prime costs of each product were as follows:

High F Great P
Direct materials $ 39.00 $ 26.40
Direct labor $ 18.60 $ 14.20
Budget factory overhead:
Engineering and Design 2,600 engineering hours $ 468,000
Quality Control 13,030 inspection hours 310,560
Machinery 33,910 machine hours 574,940
Miscellaneous Overhead 26,590 direct labor hours 185,500
Total $ 1,539,000

National's controller had been researching activity-based costing and decided to switch to it. A special study determined National's two products have the following budgeted activities:

High F Great P
Engineering and design hours 1,060 1,540
Quality control inspection hours 5,740 7,290
Machine hours 20,380 13,530
Labor hours 12,100 14,490

Using activity-based costing, total overhead per unit of the High F model is: (Rounded to the nearest cent.)
Answer #1

Calculate overhead per unit of High F model

High F
Engineering and Design 468000/2600*1060 = 190800
Quality Control 310560/13030*5740 = 136808
Machinery 574940/33910*20380 = 345540
Miscellaneous overhead 185500/26590*12100 = 84413
Total overhead 757561
Unit 11000
Overhead per unit 68.87